Using the given information to estimate the property value through Cost Approach

Cost Approach

Reproduction Costs

Assumptions

House

• Living area:2960 sf

Garage

• Living area, estimated cost of reproduction: $62.00 / sf

Appliances

• Garage, estimated cost of reproduction: $8,800

Total Reproduction

• Appliances, estimated cost of reproduction: $2,400

Depreciation

• Physical depreciation: 5% of total improvement costs

Depreciated Cost of Building

• Functional obsolescence: None

Site Value

• External obsolescence: None

Misc Site Improvements

• Site value, estimated: $40,000

Total Value

• Landscaping and other miscellaneous site improvements,

current value: $4,000.
Answer #1
Reproduction Costs
House       1,83,520 =2960*62
Garage           8,800
Appliances           2,400
Total Reproduction       1,94,720 =183520+8800+2400
Depreciation           9,736 =5% * 194720
Depreciated Cost of Building       1,84,984 =194720 - 9736
Site Value          40,000
Misc Site Improvements           4,000
Total Value       2,28,984 =184984+40000+4000